WhaT IS THE PERCENT DECREASE OF A NOTEBOOK COSTING $1,350 ON SALE FOR $1,200?
Answer by ewatrrr(24785)


Hi
Let x represent the percentage decrease
1350 - x*1350 = 1200
1-x = 1200/1350
x = 1-.889 = .111 or 11.1 % decrease